OpenStudy (anonymous):

sergio has two similar cylindrical pots, pot A and pot B. The radius of pot A is 12 inches, and the radius of pot B is 3 inches. what is the ratio of the volume of pot A to the volume of pot B?

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

volume is a cubic measure

OpenStudy (anonymous):

does this have something to do with change of demension formula?

OpenStudy (anonymous):

you have to cube the linear ratios

OpenStudy (anonymous):

dude your speaking german to me lol im so lost

OpenStudy (anonymous):

let r = smaller pot's radius and R = the bigger pot's radius then r*a = R where a is the scaling factor from small to big. then the volume of the pots involves the radius to the 3rd power: (r*a)^3 = R^3 and the scaling factor gets cubed (raised to the 3rd power)

OpenStudy (anonymous):

so plug the numbers into that and i get my answer?

OpenStudy (anonymous):

the ratio is 4:1 raise each nuber to the 3rd power... what do you get?

OpenStudy (anonymous):

12^3= 1728 and 3^3=27

OpenStudy (anonymous):

you could do it that way but you still have to reduce... it's much easier to reduce the ratio first and then raise to the 3rd power.

OpenStudy (anonymous):

12:3 = 4:1

OpenStudy (anonymous):

64 and 1

OpenStudy (anonymous):

you got it!